A girl delivering newspaper covers her route by traveling 3.00 blocks west, 4.00 blocks north, and then 6.00 blocks east. (a) What is her resultant displacement? (b) What is the total distance she travels?
1
Expert's answer
2020-10-12T11:42:29-0400

solution

for this question diagram can be shown as below



a)resultant displacement from point A to E is given by

"AE=\sqrt{AD^2+DE^2}"

"=\sqrt{4^2+3^2}"

=5 blocks

b)total distance is travelled

"d=AB+BC+CE"

"d=3+4+6=13" blocks